A man invests part of 17,000 pesos at 8% and the rest at 9%. if the annual income from the 2 investments are equal, how much has he invested at each rate?
Answer by checkley79(3341)

.08X=.09(17,000-X)
.08X=1,530-.09X
.08X+.09X=1,530
.17X=1,530
X=1,530/.17
X=9,000 AMOUNT INVESTED @ 8%.
17,000-9,000=8,000 AMOUNT INVESTED @ 9%.
PROOF:
.08*9,000=720 INTEREST.
.09*8,000=720 DITTO.